please solve me this problem (( find the area of a regular hexagon that has a side of 1 cm and an apothem approximately 8.7 cm?
Answer by edjones(8007)

A regular hexagon with a side of 1 should have an apothem of sqrt(3)/2=.866 or approximately .87 NOT 8.7
.5*.87*1=.435
.435*6=2.61 cm^2
.
Ed